9. Arondale Aeronautics has perpetual preferred stock outstanding with a par value of $100. The stock...

9.

Arondale Aeronautics has perpetual preferred stock outstanding with a par value of $100. The stock pays a quarterly dividend of $3, and its current price is $79.

  1. What is its nominal annual rate of return? Do not round your intermediate calculations. Round your answer to two decimal places.
    %
  2. What is its effective annual rate of return? Do not round your intermediate calculations. Round your answer to two decimal places.
Answer #1

a. The nominal annual rate of return is:

=Annual dividend/Current price

=3*4/79

=15.19%

b. EAR = (1+r/n)n-1

=(1+15.19%/4)^4-1

=16.08%